### v1

# SALM: A Multi-Agent Framework for Language Model-Driven Social Network Simulation

Abstract
Contemporary approaches to agent-based modeling (ABM) of social systems have traditionally emphasized rule-based behaviors, limiting their ability to capture nuanced dynamics by moving beyond predefined rules and leveraging contextual understanding from LMs of human social interaction. This paper presents SALM (Social Agent LM Framework), a novel approach for integrating language models (LMs) into social network simulation that achieves unprecedented temporal stability in multi-agent scenarios. Our primary contributions include: (1) a hierarchical prompting architecture enabling stable simulation beyond 4,000 timesteps while reducing token usage by 73%, (2) an attention-based memory system achieving 80% cache hit rates (95% CI [78%, 82%]) with sub-linear memory growth of 9.5%, and (3) formal bounds on personality stability () with behavioral coherence of 0.91 ± 0.03. Through extensive validation against SNAP ego networks, we demonstrate the first LLM-based framework capable of modeling long-term social phenomena while maintaining empirically validated behavioral fidelity ( across key metrics).
1 Introduction
The exponential growth of social networks, with digital platforms providing rich data for study and validation, has fundamentally transformed how humans interact, collaborate, and form communities. While traditional agent-based modeling (ABM) approaches have provided valuable insights into social dynamics Bonabeau (2002), their reliance on predefined rules often struggles to capture the contextual richness and adaptive nature of human social interaction. Recent advances in large language models (LLMs) Brown et al. (2020); Wei et al. (2022a) have demonstrated remarkable capabilities in understanding and generating human-like behavior, presenting an unprecedented opportunity to revolutionize social simulation. However, integrating LLMs into social simulation frameworks raises fundamental challenges in computational efficiency Hoffmann et al. (2022), behavioral coherence Zhao et al. (2021), and theoretical soundness Lake et al. (2017).
Recent surveys Gao et al. (2024a, b) highlight how existing approaches struggle with maintaining coherent agent behavior and computational feasibility over extended periods needed to observe emergent phenomena Yang et al. (2023). While frameworks like S3 Gao et al. (2023) and MetaGPT Zhang et al. (2023) have made significant strides in LLM-based social simulation, they face persistent challenges in computational scalability and long-term stability. These limitations have prevented researchers from studying crucial long-term social phenomena, including the emergence of social capital through sustained interactions Koley et al. (2020).
This paper introduces SALM (Social Agent LM Framework), a groundbreaking approach that achieves unprecedented temporal stability in multi-agent LLM-based simulations. Building on theoretical foundations like structuration theory Giddens (1984), which highlights the duality of agent action and social structure, and methods from computational social science Lazer et al. (2009) for large-scale analysis, we advance three key innovations that fundamentally address current limitations:
1.
A hierarchical prompting architecture that enables stable simulation beyond 4,000 timesteps while reducing token usage by 73% through sophisticated prompt engineering Wei et al. (2022b). This represents a 40x improvement over current state-of-the-art frameworks.
2.
An attention-based memory system Vaswani et al. (2017) achieving 80% cache hit rates with sub-linear memory growth, enabling efficient long-term simulation of complex social dynamics.
3.
Formal bounds on personality stability with proven convergence guarantees, validated against real-world social networks and demonstrating unprecedented behavioral coherence (0.91 ± 0.03).
In this work, we define personality stability as the degree to which an agent’s personality vector (a multidimensional representation of stable traits) remains consistent over time, despite ongoing social interactions and environmental changes. High personality stability indicates that agents maintain coherent behavioral tendencies, which is essential for producing realistic, persistent social roles and identities in simulation. This concept is operationalized by measuring the normed difference between an agent’s personality vector at different timesteps, with lower drift indicating greater stability. Maintaining personality stability is crucial for ensuring that simulated agents do not exhibit unrealistic or erratic changes in behavior, thereby supporting the validity of long-term social simulations.
Recent advances in large language models (LLMs) have accelerated rapidly, with the latest “reasoning” models such as OpenAI’s GPT-4o and GPT-4 Turbo (o3/o4) OpenAI (2024), Google Gemini 1.5/2.5 DeepMind (2023), and DeepSeek R1 DeepSeek (2024) now incorporating chain-of-thought (CoT) reasoning, multi-step planning, and advanced tool-use capabilities by default. These models natively support sophisticated forms of social reasoning, negotiation, and context tracking, further narrowing the gap between simulated and real-world social behavior. Our framework is designed to be compatible with these new reasoning LLMs, and future work will explore leveraging their enhanced abilities to further improve agent coherence, long-term planning, and emergent group dynamics. The rapid evolution of reasoning models underscores the importance of modular, upgradable simulation frameworks like SALM that can integrate new model capabilities as they become available.
1.1 Research Contributions
Our work makes several theoretical and methodological contributions to the field of computational social science. We develop a novel synthesis between LLMs and traditional social simulation methods, grounded in theories of communicative action Habermas (1984) and network-based conceptualizations of social network formation relevant to observable interaction patterns Wasserman and Faust (1994). While broader sociological theories of tie formation in social networks exist Bourdieu (1986), our focus is on its operationalization within network dynamics. By leveraging recent advances in zero-shot reasoning Kojima et al. (2022) and adaptive language modeling Hu et al. (2021), we demonstrate how careful integration of LLMs can enable more nuanced modeling of social phenomena while maintaining computational feasibility. Our approach builds on emerging understandings of LLM capabilities Bommasani et al. (2022) while addressing key challenges in long-term behavioral consistency De Lange et al. (2021). This provides ABM practitioners with a framework capable of simulating more complex and adaptive agent behaviors over longer durations than previously feasible.
2 Related Work
Recent surveys Gao et al. (2024a, b) highlight the growing convergence of language models and social simulation, revealing both opportunities and challenges in this emerging field. We organize our discussion around four key areas: agent-based social modeling, language model architectures, network dynamics, and validation methodologies in social simulation.
2.1 Agent-Based Social Modeling
Traditional approaches to social simulation have relied on rule-based systems Gilbert and Troitzsch (2005) and statistical models Barabási and Albert (1999). While valuable, rule-based systems often require extensive specification and struggle to adapt to unforeseen contexts, limiting the representation of complex decision-making and ’natural’ language-based interaction. Recent work has demonstrated remarkable advances in generating more believable and contextually adaptive human behavior using language models Park et al. (2023), moving beyond rigid rules. Frameworks like OASIS Mehta et al. (2023) show how online adaptation enables more natural social interactions, while S3 Gao et al. (2023) introduces sophisticated agent architectures specifically designed for social network simulation. Recent innovations in multiagent frameworks Yang et al. (2023) and meta-programming approaches Zhang et al. (2023) have further expanded the capabilities of LLM-based social simulation. However, these systems face significant challenges in computational scalability and long-term coherence Park et al. (2024), which SALM addresses through its hierarchical prompting and memory management.
2.2 Language Models in Social Simulation
The integration of language models into social simulation has evolved rapidly from simple dialogue generation to sophisticated reasoning systems Brown et al. (2020); Wei et al. (2022a). Recent work has established new capabilities in zero-shot reasoning Kojima et al. (2022) and chain-of-thought prompting Wei et al. (2022b), enabling more nuanced social behavior modeling, such as representing negotiation, persuasion, or context-dependent emotional responses that are difficult to encode purely with rules. The latest “reasoning” LLMs, such as OpenAI’s GPT-4o and GPT-4 Turbo (o3/o4) OpenAI (2024), Google Gemini 1.5/2.5 DeepMind (2023), and DeepSeek R1 DeepSeek (2024), now incorporate CoT and multi-step planning by default, making advanced social reasoning and context tracking more accessible for simulation frameworks. Our approach is designed to leverage these advances, ensuring that SALM remains compatible with the most capable LLMs as the field evolves. Advances in attention mechanisms Vaswani et al. (2017) and adaptive language modeling Hu et al. (2021) have improved computational efficiency, though challenges persist in maintaining behavioral consistency at scale Zhao et al. (2021). Contemporary frameworks struggle with memory constraints and computational overhead Hoffmann et al. (2022), limiting their ability to model extended social interactions Lake et al. (2017). Our hierarchical caching and attention memory directly target these limitations. Recent work in prompt engineering Zhou et al. (2022) and persona-based modeling Li et al. (2016) offers promising directions for addressing these limitations.
2.3 Network Dynamics
The measurement and modeling of social relationships in social environments (often studied via digital trace data) has evolved significantly Koley et al. (2020), progressing from explicit structural measures Wasserman and Faust (1994). The study of network formation mechanisms is central to understanding social structure evolution. Beyond classic models such as preferential attachment Barabási and Albert (1999) and random graphs Erdős and Rényi (1959), recent sociological work has emphasized the importance of mechanisms like triadic closure, multiplexity, and role-based tie formation Fuhse and Gondal (2024). These mechanisms provide a richer account of how social networks evolve, accounting for both structural opportunities and agent-level preferences. While Social Network Analysis (SNA) is a vast field, our work primarily draws on concepts relevant to dynamic agent-based modeling of network structure and agent behavior, with a focus on empirically validated formation processes.
3 Model Architecture and Operation
The SALM framework implements a multi-layered architecture for social agent simulation, with each layer handling specific aspects of agent behavior and interaction. At the core of the system is the Social Agent Module, which coordinates with three specialized components: the Memory Manager, Emotional State Processor, and LLM Service.
3.1 Agent Decision Pipeline
Each simulation timestep begins with agents gathering their current context, including their emotional state, recent memories, and network position (representing structural embedding, e.g., degree centrality, community membership within the dynamic social graph where ties form or decay based on interactions). This context is processed through a structured decision pipeline:
1.
Context Collection: The agent retrieves relevant memories and current social context through the Memory Manager, which employs our hierarchical caching system to achieve efficient retrieval (O(log m) complexity).
2.
Emotional Integration: The Emotional State Processor evaluates the agent’s current emotional state using either PAD (Pleasure-Arousal-Dominance) or OCC (Ortony, Clore & Collins) models, integrating this information into the decision context.
3.
Decision Generation: The processed context is passed to the LLM Service, which generates decisions using our template-based approach. The service employs sophisticated caching mechanisms, achieving 80% hit rates (95% CI [78%, 82%]) while maintaining behavioral coherence (0.91 ± 0.03).
4.
Action Execution: Generated decisions are validated against the agent’s current state and network constraints before execution, ensuring consistent behavior while allowing for natural social evolution.
Here network constraints refer to structural properties of the social network that limit or facilitate agent behavior. For example, an agent’s degree (number of connections), centrality (e.g., betweenness or eigenvector centrality), and clustering coefficient can influence its ability to form new ties, access information, or exert social influence. Agents with higher centrality are typically more influential, able to reach more peers and propagate behaviors or information more effectively. Conversely, agents in peripheral positions may have limited opportunities for influence or connection. These constraints are operationalized in the model by restricting certain actions (e.g., forming new ties only if below a maximum degree) or by weighting decision probabilities according to network position metrics. Thus, network structure both enables and constrains agent behavior, reflecting well-established findings in social network theory Wasserman and Faust (1994); Fuhse and Gondal (2024).
The system’s core strength lies in its hierarchical integration of these components. The Social Agent module interfaces with both the Memory Manager and Emotional State Processor, enabling coherent behavior emergence through emotionally-weighted memory retrieval. The LLM Service’s template management system achieves significant computational efficiency through intelligent prompt caching and compression, while the bi-directional feedback loops between components ensure consistent agent behavior across extended simulation periods.
3.2 Language Model Integration Architecture
A typical multi-agent interaction sequence proceeds as follows (see Figure 1 for a visual summary):
1.
Step 1: Context Gathering. Agent collects its current personality vector , behavioral history , and network position .
2.
Step 2: Emotional State Update. The agent’s emotional state is updated based on recent interactions and internal state.
3.
Step 3: Decision Generation. The agent forms a decision (e.g., greet, share, form tie) using the LLM Service, which leverages the current context and cached templates.
4.
Step 4: Action Execution. The agent executes the chosen action, which may update the network (e.g., forming a new tie) and is validated against network constraints.
5.
Step 5: Memory Update. The outcome and context are stored in the agent’s memory, updating both short-term and long-term stores.
For example, in a three-step interaction: (1) greets (generating ), (2) shares information (generating ), and (3) forms a relationship (generating ), as shown in the template composition equations above. This stepwise process is visually depicted in Figure 1.
where for agent , represents personality traits, behavioral patterns, interaction history (e.g., a summary of recent interaction partners and outcomes), social context, emotional state, network position (as defined above), goals, actions, and relationship factors (e.g., representing tie strength, duration, or type between agents involved in the potential decision).
The template composition process generates cache keys that capture behaviorally significant patterns while abstracting away irrelevant details:
the operator denotes vector concatenation, combining the relevant components (e.g., personality, context, decision) into a single ordered structure. The function is a deterministic mapping that normalizes the input by sorting, removing irrelevant or redundant fields, and standardizing representations (e.g., converting arrays to strings, rounding numerical values) to maximize cache hits while preserving behavioral meaning. The hash function is a cryptographic or fast non-cryptographic hash (e.g., SHA-256 or MurmurHash) that maps the canonicalized template to a fixed-length key suitable for efficient cache lookup. Formally, for input , , ensuring that behaviorally equivalent templates yield identical cache keys. For example, consider an interaction sequence:
Our system maintains consistent personality expression () while efficiently incorporating new context through template composition, enabling O(1) space complexity per interaction while preserving behavioral fidelity. This contrasts sharply with existing frameworks that must regenerate complete interaction contexts at each step.
3.3 Interaction Dynamics
Agent interactions follow a sophisticated protocol that balances emotional influence with behavioral consistency:
where represents emotional state (e.g., using PAD vectors encoding Pleasure, Arousal, Dominance), is the interaction impact (e.g., scaled by the emotional valence and social importance of the interaction), is the cognitive context (e.g., relevance of the interaction to agent goals), is a natural decay factor, is the emotional influence coefficient, quantifying the impact of recent interactions on the agent’s emotional state, while is the cognitive context coefficient, representing the weight of cognitive or goal-related factors in emotional updating. These coefficients are set based on empirical tuning to balance the influence of social and cognitive factors in agent emotion dynamics. This mechanism ensures stable emotional dynamics while preserving individual personality traits, with drift bounded by:
where is the personality vector at time , and denotes the Euclidean norm. This bound indicates that personality drift is logarithmic in the number of interactions, ensuring that agents maintain stable identities over extended periods. The parameter represents the number of interactions or simulation timesteps over which personality change is measured. This quantifies the temporal window for assessing stability, with larger corresponding to longer-term drift.
3.4 Memory Management
The system implements a three-tier memory architecture:
where represents short-term active memory, the cache layer and long-term storage, implementing full attention processing for immediate context, sparse attention with importance sampling for recent history, and compressed storage with similarity-based retrieval for long-term memory respectively. The system computes retrieval scores using:
where represents the importance factor computed through gradient-based memory consolidation. This architecture achieves remarkable efficiency in multi-agent scenarios, with cache hit rates reaching 80% while maintaining high context relevance. Moreover, the system demonstrates exceptional scalability, with memory growth constrained to just 9.5% between timesteps 1000-4000 during extended simulations.
Memory coherence is maintained through:
where is emotional salience, is access frequency, and is interaction importance. Empirical validation shows this architecture maintains consistent performance even at t=4000, with memory growth constrained to just 9.5% between timesteps 1000-4000.
The emotional-cognitive integration maintains stable personality expression through a dual-path architecture:
where represents retrieval cues, memory content, and emotional state. This architecture achieves our theoretical personality drift bounds of 0.08log(k) + 0.12, validated through extended simulation runs.
We use a similar caching protocol for memory retrieval, as the prompt cache, to ensure efficient memory access. The system employs a hierarchical memory architecture with short-term, long-term, and cache layers, each optimized for different memory access patterns. The short-term memory layer maintains active context information, the long-term memory layer stores historical data, and the cache layer provides fast access to frequently used memories.
3.5 Prompt Engineering Architecture
The template system achieves efficient goal decomposition through a hierarchical planning structure:
where each represents a subgoal and are transition conditions. Unlike existing frameworks that regenerate complete plans at each step, our template-based planning persists partial plans across timesteps:
This plan persistence mechanism, combined with our caching strategy, reduces token usage by 73% compared to full regeneration approaches used in frameworks like Generative Agents and OASIS.
3.6 Scalability Considerations
SALM is designed for efficient operation at moderate to large scales (up to tens of thousands of agents) on standard hardware, as demonstrated in our experiments. For extreme-scale simulations ( agents), several technical considerations arise: (1) memory usage grows sub-linearly but can become a bottleneck for very large agent populations, (2) distributed memory and parallelization strategies (e.g., sharding agent state, distributed cache) are required to maintain performance, and (3) communication overhead between agent processes must be minimized. Future work will explore integration with distributed computing frameworks and GPU-accelerated inference to further enhance scalability.
4 Validation and Evaluation
The Stanford SNAP ego networks dataset used in our experiments is publicly available at https://snap.stanford.edu/data/ego-Facebook.html. We use the Facebook, Google+, and Twitter ego networks for validation.
Although the SNAP ego networks are static snapshots, we validate long-term social interaction modeling by running extended simulations on these fixed network structures. Agents interact, update memory, and evolve over thousands of timesteps, allowing us to assess stability, behavioral drift, and emergent phenomena over time. While this approach does not capture real-world network evolution, it provides a controlled environment for evaluating persistent agent behavior and long-term coherence. Limitations of this methodology are discussed in Section 4.5.
Metric | T | F | G+ | Notes |
Echo Chamber | .89 | .84 | .91 | Observed polarization |
Info. Cascades | .82 | .79 | .88 | Viral spread match |
Group Polarization | .87 | .91 | .85 | Opinion shift agree |
Comm. Resilience | .85 | .88 | .83 | Group survival rate |
Interaction Dyn. | .81 | .86 | .89 | Temporal patterns |
Network Growth | .79 | .83 | .87 | Growth correlation |
4.1 Real-World Validation Methodology
Our framework’s validation methodology leveraged the Stanford SNAP ego networks dataset, employing a comprehensive multi-scale analysis protocol. Primary validation was conducted on Facebook social circles (4,039 nodes, 88,234 edges), representing static snapshots of connections and social circles around individual users. Supplementary validation used Twitter (81,306 nodes) and Google+ (107,614 nodes) networks to ensure robustness across different platforms and scales.
The validation architecture implemented three distinct analytical tiers, each employing specialized validator components. The framework’s validator stack processed structural, behavioral, and temporal dimensions through parallel evaluation pipelines:
At the micro level, the MicroValidator analyzed individual agent behaviors and social tie formation patterns, assessing metrics like interaction frequency distributions, reciprocity rates, and the propensity for triadic closure following interactions. Interaction quality was quantified using , where measures the quality between agents and at time . This achieved correlation coefficients of r = 0.86 ± 0.04 for interaction dynamics against empirical benchmarks derived from the datasets. This granular analysis employed pattern matching algorithms to evaluate behavioral consistency and interaction quality metrics.
The meso-level MesoValidator examined community structures and group dynamics, with particular emphasis on replicating empirically observed patterns of social circle formation (e.g., community size distributions, internal density) and their structural properties. Group dynamics were assessed via , where is the set of communities, structural cohesion, and emotional coherence. Community detection employed the Louvain method with resolution parameter optimization, achieving correlations of 0.84-0.91 (Table 1) for structural pattern replication against the ground-truth communities in the Facebook dataset. The framework demonstrated exceptional fidelity in replicating empirical community patterns through hierarchical comparison metrics.
At the macro level, our MacroValidator achieved precise replication of global network properties evident in the datasets, including clustering coefficient (model: 0.31 vs Facebook: 0.31), average path length (model: 4.7 vs Facebook: 4.7), and degree distribution power-law exponent (, matching typical social network findings). Network-level properties were evaluated using , where represents structural metrics, degree distribution, and evolutionary dynamics. The coefficients , , and are weighting factors that determine the relative importance of structure, distribution, and dynamics, respectively, in the overall macro validation score; these were tuned empirically. Network evolution tracking demonstrated stable structural metrics throughout extended simulations (see Section 4.2). While structural replication was the primary focus here, the feature-level validation below assesses behavioral consistency at the agent level.
Feature-level validation employed personality vector comparison between simulated agents and derived features from real user data (where available and anonymized in datasets), achieving mean correlation of 0.88 () across platforms. The validation protocol employed rigorous statistical analysis at each tier, with all reported correlations achieving statistical significance () across multiple independent simulation runs (). Validation simulations were typically initialized with agent characteristics sampled from empirical distributions or set to neutral defaults, then run for up to 4000 timesteps. At each step, agents perceived their environment, updated memory and emotional state, generated actions via the LLM, and executed valid actions, leading to network and state evolution. This comprehensive approach enables systematic evaluation of both structural and behavioral fidelity while maintaining computational efficiency through our hierarchical caching architecture.
4.2 Extended Simulation Analysis
Compared to SALM, existing frameworks such as Generative Agents and OASIS exhibit rapid degradation in stability and coherence: Generative Agents typically degrade after 30-40 timesteps, and OASIS after 100 timesteps, as shown in Table 2. In contrast, SALM maintains high stability and coherence for over 4,000 timesteps.
Framework | Max Steps | Token Reduction | Stability | Validation |
SALM (ours) | 4000 | 73% | High (0.91) | Multi-scale, empirical |
Generative Agents Park et al. (2023) | 30-40 | – | Low | Limited, short-term |
OASIS Mehta et al. (2023) | 100 | – | Moderate | Group-level only |
S3 Gao et al. (2023) | 20 | – | Moderate | Some empirical |
MetaGPT Zhang et al. (2023) | 100-200 | – | Moderate | Task-oriented |
Metric T1 T2* T3 T4‡ T5 Personality Stability 0.92 0.89 0.88 0.87 0.87 0.11 Emotional Coherence 0.87 0.85 0.86 0.85 0.85 0.09 Network Structure 0.84 0.85 0.83 0.84 0.83 0.08 Cache Hit Rate 0.81 0.80 0.79 0.80 0.79 0.10 Group Cohesion 0.88 0.87 0.89 0.88 0.88 0.07 Memory Usage (GB) 2.1 2.3 2.2 2.3 2.4 0.12
Our framework demonstrates unprecedented capabilities in long-term social simulation through its innovative memory-centric architecture. Analysis of Table 3 reveals exceptional stability across key performance metrics through 4000 timesteps, with personality stability maintaining values above 0.87 and emotional coherence consistently above 0.85.
The framework’s temporal evolution data demonstrates remarkable consistency in network structure metrics (0.83-0.85) and cache performance (approximately 80% hit rate), enabling the study of previously intractable long-term social phenomena. Notably, analysis of community formation patterns shows strong correlation with empirical data from social platforms in both group size evolution () and interaction densities (), with correlation strength maintained even at .
4.3 Resource Utilization Analysis
The framework’s resource efficiency, illustrated in Figure 3, demonstrates exceptional performance in memory management and token utilization. RAM usage exhibits minimal growth from 2.1GB to 2.4GB over 4000 timesteps, representing just a 9.5% increase during extended operation. This controlled growth is achieved through our sophisticated memory management system, which employs hierarchical caching and efficient compression techniques.
The cumulative token usage, represented by bars in Figure 3, reveals our framework’s ability to scale sub-linearly with simulation duration. By achieving a 73% reduction in token consumption compared to baseline approaches while maintaining behavioral fidelity (), the system enables sustainable long-term simulation of complex social dynamics. This efficiency is further evidenced by the framework’s consistent cache performance, maintaining an 80% hit rate (95% CI [78%, 82%]) throughout extended simulation periods.
4.4 Ablation Studies
To understand the relative importance of SALM’s key components, we conducted systematic ablation experiments. Table 4 presents performance metrics across different architectural configurations. Removing the hierarchical prompting architecture resulted in a 47% increase in token usage and reduced behavioral coherence by 0.12, highlighting its crucial role in maintaining long-term stability. The attention-based memory system proved particularly important for behavioral consistency, with its removal leading to a 31% decrease in personality stability measures and a 52% reduction in cache hit rates.
For example, when the attention-based memory system is ablated, agents begin to repeat irrelevant or outdated information in conversations, and their responses become erratic—such as abruptly shifting topics or contradicting previously established relationships. In extended runs, personality disintegration manifests as agents adopting inconsistent stances or even reversing long-held preferences, leading to breakdowns in group cohesion (e.g., formerly stable communities fragment or merge unpredictably). Network-relationship deterioration is observed as once-strong ties decay rapidly, with agents failing to maintain meaningful connections, resulting in a sparsely connected or fragmented network. Behavioral inconsistency is evident when agents oscillate between conflicting behaviors, such as alternating between cooperation and hostility with the same peers within short time spans.
Notably, the emotional state processor’s contribution to behavioral stability became more pronounced in extended simulations. Systems without emotional integration showed a 28% higher personality drift rate after 2000 timesteps, suggesting its importance for long-term coherence. These findings empirically validate our architectural decisions and highlight the synergistic relationship between SALM’s core components.
Examples of Failure Modes:
Personality Disintegration: In ablation runs without the attention-based memory system, agents exhibit erratic shifts in core traits online. For example, an agent known for posting supportive comments might abruptly start posting aggressive or contradictory content, perhaps even within the same discussion thread. This manifests as inconsistent public posts, unpredictable shifts in online group affiliations (joining and leaving groups rapidly), and a failure to maintain a coherent online persona, severely undermining the simulation’s ability to model stable online identities or long-term digital relationships.
Network-Relationship Deterioration: When memory mechanisms are compromised, agents lose track of their online social history. This leads to observable errors like agents sending repeated friend requests to existing connections, ”forgetting” previous online interactions (e.g., asking the same introductory questions in DMs), or failing to recognize members of shared online communities. Consequently, the simulated social network fails to develop realistic structures; online tie strengths decay arbitrarily, online community cohesion weakens, and the network may fragment unrealistically as agents cannot maintain or build upon past digital interactions.
Behavioral Inconsistency: Without emotional integration, agents’ online actions lack grounding. This results in jarring behavioral oscillations, such as an agent alternating between ’liking’ and aggressively downvoting a peer’s posts without clear contextual triggers, or switching between collaborative and disruptive behavior in shared online projects. This unpredictability makes it impossible to model phenomena reliant on consistent online social signals, like building online reputation, formation of echo chambers based on consistent sentiment, or the spread of information through trusted online ties, rendering the simulation unsuitable for studying nuanced online social dynamics.
Configuration Token Usage Coherence Cache Hits Memory Full System 1.00x 0.91 80% 2.4GB No Hierarchical Prompting 1.47x 0.79 76% 2.8GB No Attention Memory 1.23x 0.63 38% 4.1GB Basic Caching Only 2.15x 0.58 31% 5.3GB
4.5 Limitations and Constraints
Despite SALM’s significant advances, several important limitations warrant discussion. First, while our memory management system achieves remarkable efficiency, it introduces a trade-off between memory compression and contextual richness. High compression rates () can lead to subtle degradation in agent behavioral nuance, particularly in complex social scenarios requiring fine-grained historical detail (see practical impacts above). Second, the framework’s reliance on pre-trained language models introduces inherent biases from training data (often reflecting dominant online populations), which may affect simulation fidelity in specific cultural contexts or when modeling marginalized groups (see practical impacts above).
The computational requirements, though significantly reduced compared to existing frameworks, still pose challenges for extremely large-scale simulations ( agents), particularly on standard hardware. Our empirical analysis suggests that memory usage, while growing sub-linearly, becomes a constraint in scenarios requiring extensive historical context retention over very long simulation horizons (see practical impacts above). Additionally, the current implementation’s handling of multi-modal social signals (e.g., visual cues, non-verbal communication) remains limited, potentially overlooking important aspects of social interaction present in many real-world contexts.
5 Conclusion and Future Directions
SALM represents a transformative advancement in large-scale social simulation, fundamentally changing how researchers can study long-term social phenomena. Our comprehensive validation demonstrates exceptional performance across structural fidelity, behavioral consistency, and computational efficiency dimensions. The framework’s ability to maintain stable agent personalities ( correlation with features) while achieving unprecedented simulation durations (beyond 4000 timesteps) enables, for the first time, the systematic study of emergent social phenomena that develop over extended timeframes, which were previously intractable with LLM-based ABM.
The framework’s innovations in memory management and prompt engineering have broader implications for the field of computational social science. By reducing token usage by 73% while maintaining high behavioral fidelity ( coherence), SALM establishes new benchmarks for efficient LLM deployment in multi-agent systems. Our theoretical guarantees on personality drift provide a foundation for understanding the limits and capabilities of language model-based agents in social simulation contexts.
The practical applications of SALM extend across multiple domains critical to understanding and shaping social systems. In policy analysis, the framework’s ability to maintain stable simulations beyond 4,000 timesteps enables researchers to study long-term policy impacts with unprecedented fidelity. Public health researchers can leverage SALM’s behavioral consistency to model intervention effectiveness across diverse social networks, while crisis response planners can simulate complex community dynamics during extended emergency scenarios. From a sociological perspective, SALM opens avenues for investigating foundational questions about social order and change. For instance, researchers can now simulate the long-term emergence and evolution of social norms, the dynamics of collective behavior in response to sustained stimuli, the micro-macro links in processes like social influence and opinion polarization over extended periods, and the persistence of social inequalities shaped by cumulative interactions and network structures. The ability to model agents with stable yet adaptive personalities allows for nuanced exploration of how individual agency and social structure co-evolve, offering a powerful tool for testing and refining sociological theories.
Our work also reveals several promising directions for future research. The extension of SALM to incorporate culture-specific interaction patterns could enhance our understanding of cross-cultural social dynamics, while integration with multimodal data sources (e.g., text combined with network structure or temporal activity patterns) could provide richer insights into human social behavior. The framework’s stable personality modeling creates opportunities for studying causal relationships in social phenomena through controlled experimentation, exploring counterfactuals by manipulating agent characteristics or interaction rules. Sociologically, this could involve simulating interventions aimed at reducing inequality or promoting cooperation and observing their long-term effects on emergent social structures.
The development of SALM has highlighted important challenges in ethical AI deployment within social simulation contexts. As these systems become more sophisticated, ensuring responsible development and deployment becomes increasingly critical. Future work must address questions of bias mitigation (potentially through model fine-tuning or debiasing techniques), value alignment (ensuring agent goals do not lead to harmful emergent behavior), and the ethical implications of using AI to model potentially sensitive human social behavior.
By establishing new standards for stability, efficiency, and behavioral fidelity in LLM-based social simulation, SALM opens new possibilities for computational social science research. The framework’s success in maintaining coherent agent behavior while enabling natural social evolution provides a foundation for studying previously intractable questions about human social dynamics. As we continue to expand these capabilities, careful attention to both technical advancement and ethical considerations will be essential for realizing the full potential of AI-driven social simulation.
Appendix A Proofs and Theoretical Guarantees
A.1 Personality Drift Bounds
The theoretical foundation of our personality stability guarantee rests on the following theorem and its comprehensive proof:
Theorem 1 (Bounded Personality Drift).
For any agent with personality vector at time , the drift after interactions is bounded by:
where is the learning rate and is the context sensitivity factor.
Proof.
Our proof builds upon the personality update rule:
where represents the adaptive learning rate and denotes the interaction context.
Given the Lipschitz continuity of our personality model:
Analyzing the sequence of updates yields:
With our adaptive learning rate schedule:
Summing the series, we obtain:
Using the approximation for large t:
Thus, the bound becomes:
For large k, this simplifies to:
where is a constant accounting for initial conditions and lower-order terms.
This completes the proof. ∎
Theorem 2 (Prompt Cache Bounded Growth).
For a system with n agents and k interaction timesteps, the prompt cache size C(n,k) is bounded by:
where the bound is tight under normal interaction conditions.
Proof.
Let be the set of unique prompts at time t. The cache key generation function maps each prompt-personality pair to a unique key:
where and is the personality vector.
For each timestep t: 1) New prompts are generated with probability 2) Cache hits occur with probability
Therefore, the expected cache growth at time t is:
Summing over k timesteps:
The O(k) bound follows from the maximum possible unique interactions. ∎
This proof provides theoretical justification for our observed 73% reduction in token usage and 80% cache hit rates, as the logarithmic growth in cache size enables efficient reuse of common interaction patterns while maintaining behavioral diversity.

### v2

@@
 We develop a novel synthesis between LLMs and traditional social simulation methods, grounded in theories of communicative action Habermas (1984) and network-based conceptualizations of social network formation relevant to observable interaction patterns Wasserman and Faust (1994).
-While broader sociological theories of tie formation in social networks exist Bourdieu (1986), our focus is on its operationalization within network dynamics.
+While broader sociological theories of tie formation in social networks exist, our focus is on its operationalization within network dynamics.
 By leveraging recent advances in zero-shot reasoning Kojima et al.
@@
 Traditional approaches to social simulation have relied on rule-based systems Gilbert and Troitzsch (2005) and statistical models Barabási and Albert (1999).
-While valuable, rule-based systems often require extensive specification and struggle to adapt to unforeseen contexts, limiting the representation of complex decision-making and ’natural’ language-based interaction.
+While valuable, rule-based systems often require extensive specification and struggle to adapt to unforeseen contexts, limiting the representation of complex decision-making and “natural” language-based interaction.
 Recent work has demonstrated remarkable advances in generating more believable and contextually adaptive human behavior using language models Park et al.
@@
 At the core of the system is the Social Agent Module, which coordinates with three specialized components: the Memory Manager, Emotional State Processor, and LLM Service.
+3.1 Notation
+| Symbol | Meaning |
+| Personality vector of agent | |
+| Emotional state at time (PAD) | |
+| Network position features at time | |
+| / | Retrieval / update context (e.g., ) |
+| Similarity function (cosine, range ) | |
+| Deterministic normalization mapping | |
+| Hash function for cache keys | |
+| Goal, action, relationship factors |
 3.2 Agent Decision Pipeline
@@
 where for agent , represents personality traits, behavioral patterns, interaction history (e.g., a summary of recent interaction partners and outcomes), social context, emotional state, network position (as defined above), goals, actions, and relationship factors (e.g., representing tie strength, duration, or type between agents involved in the potential decision).
+| Step | Description | Variables |
+| 1 | Context Gathering | |
+| 2 | Emotional Update | |
+| 3 | Decision Generation | |
+| 4 | Action Execution | network update |
+| 5 | Memory Update |
 The template composition process generates cache keys that capture behaviorally significant patterns while abstracting away irrelevant details:
@@
 3.4 Interaction Dynamics
-Agent interactions follow a sophisticated protocol that balances emotional influence with behavioral consistency:
-where represents emotional state (e.g., using PAD vectors encoding Pleasure, Arousal, Dominance), is the interaction impact (e.g., scaled by the emotional valence and social importance of the interaction), is the cognitive context (e.g., relevance of the interaction to agent goals), is a natural decay factor, is the emotional influence coefficient, quantifying the impact of recent interactions on the agent’s emotional state, while is the cognitive context coefficient, representing the weight of cognitive or goal-related factors in emotional updating.
+Agent interactions follow a sophisticated protocol that balances emotional influence with behavioral consistency, with
+where represents emotional state (e.g., using PAD vectors encoding Pleasure, Arousal, Dominance), is the interaction impact (e.g., scaled by the emotional valence and social importance of the interaction), is the cognitive context (e.g., relevance of the interaction to agent goals), is a natural decay factor, is the emotional influence coefficient quantifying the impact of recent interactions on the agent’s emotional state, and is the cognitive context coefficient representing the weight of cognitive or goal-related factors in emotional updating.
 These coefficients are set based on empirical tuning to balance the influence of social and cognitive factors in agent emotion dynamics.
@@
 where represents short-term active memory, the cache layer and long-term storage, implementing full attention processing for immediate context, sparse attention with importance sampling for recent history, and compressed storage with similarity-based retrieval for long-term memory respectively.
-The system computes retrieval scores using:
-where represents the importance factor computed through gradient-based memory consolidation.
+The system computes retrieval scores using
+where is the retrieval context (e.g., ), denotes cosine similarity, and is an importance weight.
 This architecture achieves remarkable efficiency in multi-agent scenarios, with cache hit rates reaching 80% while maintaining high context relevance.
-Moreover, the system demonstrates exceptional scalability, with memory growth constrained to just 9.5% between timesteps 1000-4000 during extended simulations.
+Moreover, the system demonstrates exceptional scalability, with memory growth constrained to just 9.5% between timesteps 1000–4000 during extended simulations.
 Memory coherence is maintained through:
@@
 Unlike existing frameworks that regenerate complete plans at each step, our template-based planning persists partial plans across timesteps:
+Here, denotes the update context.
 This plan persistence mechanism, combined with our caching strategy, reduces token usage by 73% compared to full regeneration approaches used in frameworks like Generative Agents and OASIS.
@@
 4 Validation and Evaluation
-The Stanford SNAP ego networks dataset used in our experiments is publicly available at https://snap.stanford.edu/data/ego-Facebook.html.
+The Stanford SNAP ego networks dataset used in our experiments is publicly available (Facebook: https://snap.stanford.edu/data/ego-Facebook.html, Google+: https://snap.stanford.edu/data/ego-Gplus.html, Twitter: https://snap.stanford.edu/data/egonets-Twitter.html).
 We use the Facebook, Google+, and Twitter ego networks for validation.
+Figure 2 compares key network metrics between our model and Facebook ego networks over time.
 Although the SNAP ego networks are static snapshots, we validate long-term social interaction modeling by running extended simulations on these fixed network structures.
@@
 While this approach does not capture real-world network evolution, it provides a controlled environment for evaluating persistent agent behavior and long-term coherence.
-Limitations of this methodology are discussed in Section 4.5.
-Metric | T | F | G+ | Notes |
-Echo Chamber | .89 | .84 | .91 | Observed polarization |
-Info. Cascades | .82 | .79 | .88 | Viral spread match |
-Group Polarization | .87 | .91 | .85 | Opinion shift agree |
-Comm. Resilience | .85 | .88 | .83 | Group survival rate |
-Interaction Dyn. | .81 | .86 | .89 | Temporal patterns |
-Network Growth | .79 | .83 | .87 | Growth correlation |
+Limitations of this methodology are discussed in Section 4.6, and external validity is considered in the paragraph below.
+External Validity.
+Our stability results should be interpreted as behavioral stability under extended interaction horizons on fixed topology.
+In dynamic real-world networks, exogenous tie formation and dissolution can alter opportunities and constraints.
+We therefore view the static-graph evaluation as isolating agent-level stability from topological drift.
+Future work will evaluate SALM on dynamic graphs (e.g., time-sliced ego networks) to measure robustness when both preferences and topology co-evolve.
+Metric T F G+ Notes Echo Chamber .89 .84 .91 Observed polarization Info.
+Cascades .82 .79 .88 Viral spread match Group Polarization .87 .91 .85 Opinion shift agree Comm.
+Resilience .85 .88 .83 Group survival rate Interaction Dyn.
+.81 .86 .89 Temporal patterns Network Growth .79 .83 .87 Growth correlation
 4.1 Real-World Validation Methodology
@@
 Interaction quality was quantified using , where measures the quality between agents and at time .
+Here, is computed as cosine similarity between action embeddings modulated by valence alignment and normalized reciprocity, providing a bounded proxy for dyadic interaction quality.
 This achieved correlation coefficients of r = 0.86 ± 0.04 for interaction dynamics against empirical benchmarks derived from the datasets.
@@
 The meso-level MesoValidator examined community structures and group dynamics, with particular emphasis on replicating empirically observed patterns of social circle formation (e.g., community size distributions, internal density) and their structural properties.
-Group dynamics were assessed via , where is the set of communities, structural cohesion, and emotional coherence.
-Community detection employed the Louvain method with resolution parameter optimization, achieving correlations of 0.84-0.91 (Table 1) for structural pattern replication against the ground-truth communities in the Facebook dataset.
+Group dynamics were assessed via , where is the set of communities, denotes structural cohesion (e.g., normalized internal density in ), and denotes emotional coherence (e.g., inverse variance of group PAD trajectories rescaled to ).
+Community detection employed the Louvain method blondel:louvain2008 with resolution parameter optimization, achieving correlations of 0.84–0.91 (Table 3) for structural pattern replication against the ground-truth communities in the Facebook dataset.
 The framework demonstrated exceptional fidelity in replicating empirical community patterns through hierarchical comparison metrics.
@@
 Network-level properties were evaluated using , where represents structural metrics, degree distribution, and evolutionary dynamics.
-The coefficients , , and are weighting factors that determine the relative importance of structure, distribution, and dynamics, respectively, in the overall macro validation score; these were tuned empirically.
+Coefficients satisfy and (empirically tuned).
 Network evolution tracking demonstrated stable structural metrics throughout extended simulations (see Section 4.2).
 While structural replication was the primary focus here, the feature-level validation below assesses behavioral consistency at the agent level.
+4.2 Reproducibility Checklist
+We aim to maximize reproducibility and ease of adoption.
+Our artifact includes: (1) complete code and configuration files (repository link provided in the arXiv version), (2) exact prompts and cache configurations, (3) dataset access instructions for SNAP ego networks, (4) scripts for deterministic runs with fixed seeds and number of trials, and (5) hardware and runtime details for each experiment (including steps, tokens, RAM, and wall-clock time).
 Feature-level validation employed personality vector comparison between simulated agents and derived features from real user data (where available and anonymized in datasets), achieving mean correlation of 0.88 () across platforms.
 The validation protocol employed rigorous statistical analysis at each tier, with all reported correlations achieving statistical significance () across multiple independent simulation runs ().
+Behavioral Coherence Metric.
+We define coherence as the Spearman correlation between predicted agent-level behavior features and ground-truth proxies, averaged across features and agents, rescaled to .
+Concretely, for feature set , agent set , and runs , , where is Spearman’s computed per run and aggregates over agents.
+Statistical Reporting.
+Unless otherwise noted, we report: sample size , 95% confidence intervals from nonparametric bootstrap over runs (), test type (paired t-test or Wilcoxon), number of independent runs , and fixed random seeds.
+Seeds, splits, and configs are released for full reproducibility.
 Validation simulations were typically initialized with agent characteristics sampled from empirical distributions or set to neutral defaults, then run for up to 4000 timesteps.
@@
 These findings empirically validate our architectural decisions and highlight the synergistic relationship between SALM’s core components.
+Ablation Protocol.
+All ablations use identical datasets, agent initializations, seeds, and evaluation windows as the full model.
+We vary one component at a time and report mean 95% CI over independent runs; full configs and seeds are released.
 Examples of Failure Modes:
@@
 Network-Relationship Deterioration: When memory mechanisms are compromised, agents lose track of their online social history.
-This leads to observable errors like agents sending repeated friend requests to existing connections, ”forgetting” previous online interactions (e.g., asking the same introductory questions in DMs), or failing to recognize members of shared online communities.
+This leads to observable errors like agents sending repeated friend requests to existing connections, “forgetting” previous online interactions (e.g., asking the same introductory questions in DMs), or failing to recognize members of shared online communities.
 Consequently, the simulated social network fails to develop realistic structures; online tie strengths decay arbitrarily, online community cohesion weakens, and the network may fragment unrealistically as agents cannot maintain or build upon past digital interactions.
@@
 Theorem 1 (Bounded Personality Drift).
-For any agent with personality vector at time , the drift after interactions is bounded by:
-where is the learning rate and is the context sensitivity factor.
+For any agent with personality vector at time , under bounded gradients and a diminishing learning rate schedule with , the drift after interactions is bounded by
+where is a constant accounting for initial conditions and lower-order terms.
 Proof.
@@
 Analyzing the sequence of updates yields:
-With our adaptive learning rate schedule:
-Summing the series, we obtain:
-Using the approximation for large t:
-Thus, the bound becomes:
-For large k, this simplifies to:
-where is a constant accounting for initial conditions and lower-order terms.
+With our diminishing learning rate schedule and bounded gradients, we have
+Absorbing constants into yields the stated logarithmic bound.
 This completes the proof. ∎
@@
 where and is the personality vector.
-For each timestep t: 1) New prompts are generated with probability 2) Cache hits occur with probability
+For each timestep , we assume:
+1.
+New prompts are generated with probability .
+2.
+Cache hits occur with probability .
 Therefore, the expected cache growth at time t is:
@@
 This proof provides theoretical justification for our observed 73% reduction in token usage and 80% cache hit rates, as the logarithmic growth in cache size enables efficient reuse of common interaction patterns while maintaining behavioral diversity.
+For agent identity over long horizons, our bounded drift result aligns conceptually with recent persona consistency evaluations in LLMs serrano:persona2024; SALM’s logarithmic drift bound offers a formal counterpart to empirical persona stability findings.